OpenStudy (anonymous):

On a number line, point A has coordinate -4, and point B has coordinate 5. Find the coordinate of the point two thirds of the way from A to B.

OpenStudy (anonymous):

find the distance between A and B 5 - (-4) = 5 + 4 = 9 find two thirds of 9 = 6 add this on to -4 to get 2